anna is three times as old mica now,but in three years time she will only be twice as old. how old is anna and mica now?
Answer by math_helper(2461)


A good problem to solve in your head.
The algebraic formulation is:
A = Anna's age now
M = Mica's age now
A = 3M (1)
A + 3 = 2(M+3) (2)
Substitute 3M for A and expand Right Hand Side (RHS):
3M + 3 = 2M + 6
M = 3 --> Mica is 3, Anna is 3*3 = 9